what volume does 81.2 of CH4 occupy at 5.5° celcius and 31 kPa


1
Expert's answer
2021-04-25T06:15:14-0400

P = 31 kPa

V = unknown

R = 8.314 L kPa mol-1 K-1

T = 5.5°C = 278.5 K

m(CH4) = 81.2 g


Solution:

The molar mass of CH4 is 16 g/mol.

Hence,

Moles of CH4 = (81.2 g CH4) × (1 mol CH4 / 16 g CH4) = 5.075 mol

n(CH4) = 5.075 mol


In order to solve this problem we would use the Ideal Gas Law formula:

PV = nRT

Divide to isolate V:

V = nRT / P

Plug in the numbers and solve for V:

V = (5.075 mol × 8.314 L kPa mol-1 K-1 × 278.5 K) / (31 kPa) = 379.06 L = 379 L

V = 379 L


Answer: V = 379 L